### v1

# Certifying Anosov representations

Abstract.
By providing new finite criteria which certify that a finitely generated subgroup of $\operatorname{SL}(d,\operatorname{\mathbb{R}})$ or $\operatorname{SL}(d,\operatorname{\mathbb{C}})$ is projective Anosov, we obtain a practical algorithm to verify the Anosov condition. We demonstrate on a surface group of genus 2 in $\mathrm{SL}(3,\mathbb{R})$ by verifying the criteria for all words of length 8. The previous version required checking all words of length $2$ million.
1. Introduction
In general, it is a hard problem to determine whether a finite subset of $\operatorname{SL}(d,\operatorname{\mathbb{R}})$ generates a discrete subgroup. Aside from the special case of $\operatorname{SL}(2,\operatorname{\mathbb{R}})$ [GM91], the discreteness property cannot be decided by an algorithm [Kap16]. Nonetheless, it remains possible to verify stronger properties than discreteness via certain geometric algorithms [Kap23]. This paper concerns a numerically stable algorithm that can verify the Anosov property, which is stronger than discreteness, in finite time.
Anosov subgroups, introduced by Labourie [Lab06] and Guichard-Wienhard [GW12], provide a rich yet tractable source of examples of infinite covolume discrete subgroups of higher rank Lie groups. Their importance is underscored by their central role in higher Teichmüller theory [Hit92, FG06, GW18] as well as the rich examples of dynamical systems [CZZ24, ELO23, Sam24] and geometric structures on manifolds [KLP18a, DGK23, Kas18] that they provide. Anosov subgroups generalize convex cocompact subgroups of isometries of hyperbolic space to higher rank. In particular, they can be characterized in terms of the coarse geometry of their action on the associated symmetric space [KLP17, KLP18b, BPS19].
In the present article, we present the first practical algorithm which certifies that a finitely generated linear group is projective Anosov. The core of the algorithm, and main result of this paper, is Theorem 4.2, which establishes new finite criteria for a sequence in the symmetric space to have coarsely linear singular value gaps. We demonstrate the practicality of the algorithm by verifying the criteria on an example of a surface group in $\operatorname{SL}(3,\operatorname{\mathbb{R}})$ by checking words of length $8$.
The approach is based on a coarse geometric characterization of Anosov subgroups due Kapovich-Leeb-Porti [KLP14, KLP23]. They proved a local-to-global property for Morse quasigeodesics, and described an algorithm to certify the Anosov property of a finitely generated subgroup of a semisimple Lie group. If a subgroup is Anosov, their algorithm will stop and certify so in finite time; otherwise, if the subgroup is not Anosov, the algorithm will run forever. The author made their algorithm effective [Rie21] by supplying their arguments with explicit estimates. While effective, that version of the algorithm was impractical, requiring the user to verify a condition on words of length $2$ million even on a simple example of an Anosov surface group in $\operatorname{SL}(3,\operatorname{\mathbb{R}})$.
The bulk of the improvement is due to Lemma 3.1, which provides a formula relating $\zeta$-angles (Definition 3.5) and distances to parallel sets (§2.7) of the relevant type. Such a formula cannot hold for parallel sets in general (Remark 3.2). This is why the present paper only directly deals with projective Anosov subgroups of $\operatorname{SL}(d,\operatorname{\mathbb{R}})$ or $\operatorname{SL}(d,\operatorname{\mathbb{C}})$, rather than general $\Theta$-Anosov subgroups of an arbitrary semisimple Lie group $G$. Fortunately, verifying the $\Theta$-Anosov property of such a subgroup reduces to the projective Anosov case in a straightforward way [GW12, Section 4]. We mention that a different algorithm to verify the $\Theta$-Anosov property was recently obtained by the author and Davalo [DR24]. It is based on Dirichlet domains with respect to Finsler metrics, and guaranteed to eventually terminate for $\Theta$-Anosov subgroups in certain cases, e.g. $n$-Anosov subgroups of $\operatorname{Sp}(2n,\operatorname{\mathbb{R}})$.
The local-to-global principle of Kapovich-Leeb-Porti relies on a Theorem which guarantees that sufficiently straight and spaced sequences are Morse quasigeodesics. Our Theorem 4.2 is a similar but slightly different statement that guarantees a sequence is $d_{\alpha}$-undistorted (Definition 4.1), which means that the first singular value gap grows coarsely linearly. Besides the statement, there are also some technical differences with the proof here compared with their proof and that of [Rie21]. For example, Kapovich-Leeb-Porti work with an $\iota$-invariant model simplex (where $\iota$ is the “opposition involution”) and we crucially drop that assumption here. Compared to the proof in [Rie21] we make use of further auxiliary parameters and incorporate new estimates (Lemma 2.2 and Lemma 3.6) and the key angle-to-distance formula, Lemma 3.1, mentioned above.
Applying the local-to-global principle amounts to calculating various geometric quantities in the associated symmetric space. An implementation by the author is available at [Rie24], and a faster implementation by Teddy Weisman is available at [Wei24]. Both implementations are in Python. KBMAG [HGT23] is required to produce an automoton and enumerate all geodesic words of length $8$. Strictly speaking, neither computation is rigorous, in the sense that there is no guarantee on the numerical precision in the calculation; however, the results presented here are compatible with the computation in [Rie23] based on hyperbolic geometry, so are expected to be accurate. An implementation with numerical guarantees is necessary to use these techniques to rigorously prove the Anosov property. A generalized implementation, which can accept approximate generating sets and rigorously guarantee the Anosov condition, is an appealing avenue for future work.
Acknowledgements. I remain grateful to my PhD advisor, Jeff Danciger, who first suggested to me the problem of making the Kapovich-Leeb-Porti algorithm effective.
2. Symmetric space reminders
For necessary background on symmetric spaces we refer to [Hel79, Ebe96].
2.1. The model and metric
Let $\operatorname{\mathbb{K}}=\operatorname{\mathbb{R}}$ or $\operatorname{\mathbb{C}}$ and let $\operatorname{\mathbb{X}}$ denote the symmetric space associated to $\operatorname{SL}(d,\operatorname{\mathbb{K}})$. Concretely, a model for $\operatorname{\mathbb{X}}$ is given by
$\operatorname{\mathbb{X}}=\{X\in\operatorname{Herm}(d,\operatorname{\mathbb{K}}):X\gg 0,\det(X)=1\}$
where $\operatorname{Herm}(d,\operatorname{\mathbb{K}})$ denotes Hermitian $d\times d$ matrices with entries in $\mathbb{K}$ and $X\gg 0$ means that $X$ is positive definite. There is a natural action of $\operatorname{SL}(d,\operatorname{\mathbb{K}})$ on $\operatorname{\mathbb{X}}$ given by $g.X=gXg^{\dagger}$. We denote the stabilizer of the identity matrix $I_{d}$ by $\operatorname{\mathrm{SU}}(d,\operatorname{\mathbb{K}})$, which is simply $\operatorname{SO}(d)$ if $\operatorname{\mathbb{K}}=\operatorname{\mathbb{R}}$ and $\operatorname{\mathrm{SU}}(d)$ if $\operatorname{\mathbb{K}}=\operatorname{\mathbb{C}}$.
There is a unique $\operatorname{SL}(d,\operatorname{\mathbb{K}})$-invariant Riemannian metric on $\operatorname{\mathbb{X}}$ up to scale. With such a metric, $\operatorname{\mathbb{X}}$ becomes a Riemannian symmetric space: for each $p\in\operatorname{\mathbb{X}}$ there exists a (unique) involutive isometry $S_{p}\colon\operatorname{\mathbb{X}}\to\operatorname{\mathbb{X}}$ with $p$ as an isolated fixed point. Moreover $\operatorname{\mathbb{X}}$ is a symmetric space of non-compact type and in particular $\operatorname{\mathbb{X}}$ is a Hadamard manifold.
In this paper we use the Riemannian metric:
$\forall X,Y\in T_{p}\operatorname{\mathbb{X}}\subset\operatorname{Herm}(d,\operatorname{\mathbb{K}}),\quad\left\langle{X},{Y}\right\rangle_{p}=\frac{1}{2}\operatorname{Tr}(p^{-1}Xp^{-1}Y).$
The specific choice of metric will simplify some of the formulas to follow. For this metric, when $d\geq 3$, the sectional curvature of $\operatorname{\mathbb{X}}$ has image $[-1,0]$, and for $d=2$, $\operatorname{\mathbb{X}}$ has constant sectional curvature $-1$.
2.2. Geodesics and vector-valued distance
If $c\colon\operatorname{\mathbb{R}}\to\operatorname{\mathbb{X}}$ is a geodesic, then there is a unique $1$-parameter subgroup $t\mapsto\exp(tX)$ in $\operatorname{SL}(d,\operatorname{\mathbb{K}})$ such that $c(t)=\exp(tX).c(0)$. We now recall the well-known fact that every geodesic in a symmetric space can be put into a standard position. Let $\mathfrak{a}$ denote the set of real diagonal $d\times d$ matrices of trace 00, and let $\mathfrak{a}^{+}$ denote the subset of $\mathfrak{a}$ whose diagonal entries are non-increasing.
Theorem 2.1.
If $c$ is a geodesic with $c(0)=I_{d}$ then there exist $k,k^{\prime}\in\operatorname{\mathrm{SU}}(d,\operatorname{\mathbb{K}})$ and $A\in\mathfrak{a}^{+}$ such that $c(t)=k\exp(tA)k^{\prime}.I_{d}$. The element $A\in\mathfrak{a}^{+}$ is uniquely determined.
This leads to a vector-valued invariant for tangent vectors and a vector-valued distance for pairs of points. Indeed, if $v\in T_{p}\operatorname{\mathbb{X}}$, then let $c$ be the geodesic with $c^{\prime}(0)=v$, and set $\vec{d}(v)=A\in\mathfrak{a}^{+}$. Similarly, for a pair of points $p,q\in\operatorname{\mathbb{X}}$ we let $\vec{d}(p,q)$ denote the unique $A\in\mathfrak{a}^{+}$ associated to the geodesic segment $pq$.
2.3. Scale of the metric
The map $\mathfrak{a}\to\mathbb{X}$ given by $A\mapsto\exp(A).I_{d}$ becomes an isometry when $\mathfrak{a}$ is endowed with the inner product
$A,B\in\mathfrak{a}\mapsto 2\operatorname{Tr}(AB),$
because the derivative of the orbit map at $I_{d}$ restricted to symmetric matrices is multiplication by $2$. Note that the Killing form of $\mathfrak{sl}(d,\operatorname{\mathbb{K}})$ is given by
$X,Y\in\mathfrak{sl}(d,\operatorname{\mathbb{K}})\mapsto 2d\operatorname{Tr}(XY),$
so the metric we use in this paper is $\frac{1}{d}$ times the Riemannian metric induced by the Killing form. As a result, the present paper has some slightly different formulas compared to [Rie21]. Specifically, each appearance of $\kappa_{0}$ there is replaced with $1$ here.
2.4. The visual boundary
The visual boundary of $\operatorname{\mathbb{X}}$, denoted $\partial_{\rm vis}\operatorname{\mathbb{X}}$, is the set of unit-speed geodesic rays up to asymptotic equivalence. For any $p\in\operatorname{\mathbb{X}}$, the exponential map yields a homeomorphism $S(T_{p}\operatorname{\mathbb{X}})\to\partial_{\rm vis}\operatorname{\mathbb{X}}$ with respect to the visual topology on $\partial_{\rm vis}\operatorname{\mathbb{X}}$. Any isometry of $\operatorname{\mathbb{X}}$ extends to a homemorphism of $\partial_{\rm vis}\operatorname{\mathbb{X}}$. A fundamental domain for the action of $\operatorname{SL}(d,\operatorname{\mathbb{K}})$ on $\partial_{\rm vis}\operatorname{\mathbb{X}}$ is given by the set of unit vectors in $\mathfrak{a}^{+}$, which we denote by $\sigma_{mod}$ (called the model (spherical) Weyl chamber). Each $\operatorname{SL}(d,\operatorname{\mathbb{K}})$-orbit in $\partial_{\rm vis}\operatorname{\mathbb{X}}$ meets $\sigma_{mod}$ exactly once, so there is an induced map $\partial_{\rm vis}\operatorname{\mathbb{X}}\to\sigma_{mod}$. The image of an ideal point under this map is called its type.
2.5. Projective space
This paper concerns projective Anosov subgroups, so projective space and its dual will play a distinguished role. We must first describe how those spaces are embedded naturally in the visual boundary.
Let
$Z=\frac{1}{\sqrt{2d(d-1)}}\begin{bmatrix}d-1&&&&\\ &-1&&&\\ &&-1&&\\ &&&\ddots&\\ &&&&-1\\ \end{bmatrix}\in\mathfrak{a}^{+}$
and consider the unit-speed geodesic ray $c_{Z}(t)\coloneqq\exp(tZ).I_{d}$. The orbit $G.[c_{Z}]\subset\partial_{\rm vis}\operatorname{\mathbb{X}}$ is a copy of the projective space $\mathbb{KP}^{d-1}$ with equivariant diffeomorphism given by $g[c_{Z}]\mapsto g.[e_{1}]$, where $e_{1}$ spans the $(d-1)$-eigenspace of $Z$. Simililarly $G.[c_{-Z}]\subset\partial_{\rm vis}\operatorname{\mathbb{X}}$ is a copy of the dual projective space $(\mathbb{KP}^{d-1})^{\ast}$. For the rest of the paper, we will abuse notation by taking this identification to be implicit. In particular, we will frequently write $\measuredangle_{q}(\hat{\tau},\tau)$ for the Riemannian angle at $q$ between the ideal points $\tau\in\mathbb{KP}^{d-1}$ and $\hat{\tau}\in(\mathbb{KP}^{d-1})^{\ast}$.
For consistency with the notation of [KLP14, Rie21], we will use $\zeta$ to denote the type of $[c_{Z}]$ and $\iota\zeta$ to denote the type of $[c_{\iota Z}]$, where
$\iota Z=\frac{1}{\sqrt{2d(d-1)}}\begin{bmatrix}1&&&&\\ &1&&&\\ &&1&&\\ &&&\ddots&\\ &&&&1-d\\ \end{bmatrix}\in\mathfrak{a}^{+}.$
2.6. Regular directions and root pseudometric
A geodesic segment $pq$ or tangent vector $v$ is $\zeta$-regular if its vector-valued invariant $A=\operatorname{Diag}(a_{1},\dots,a_{d})$ has $a_{1}-a_{2}>0$. This occurs if and only if $qp$ (resp. $-v$) is $\iota\zeta$-regular, i.e. their vector-valued invariant is $B$ with $b_{d-1}-b_{d}>0$. For a $\zeta$-regular ideal point $\xi\in\partial_{\rm vis}\operatorname{\mathbb{X}}$, there exists a unique $\tau=\zeta(\xi)\in\mathbb{KP}^{d-1}$ such that every Weyl chamber containing $\xi$ also contains $\tau$, and likewise for $\iota\zeta$-regular points.
We will also be interested in a quantified version of regularity. The (first simple) root pseudometric $d_{\alpha}(x,y)$ is given by $a_{1}-a_{2}$ when its vector-valued distance is $A=\operatorname{Diag}(a_{1},\dots,a_{d})$. We record $\zeta_{0}\coloneq d_{\alpha}(Z)=\sqrt{\frac{d}{2(d-1)}}\geq\frac{1}{\sqrt{2}}$; this constant will appear in Theorem 4.2. We note that, if $\sigma_{i}(x)$ denotes the $i$th singular value of $x\in\operatorname{\mathbb{X}}$, then the root pseudodistance to the basepoint satisfies
$d_{\alpha}(I_{d},x)=\frac{1}{2}\log\left(\frac{\sigma_{1}(x)}{\sigma_{2}(x)}\right).$
We will make use of a comparison between the root pseudometric and certain Busemann functions. The assignment of Busemann functions to geodesic rays descends to an identification of the visual boundary with Busemann functions modulo constant functions. We consider the partial flag manifold $\mathcal{F}$ of line-hyperplane pairs, embedded into the visual boundary so that every type corresponds to a coroot. Alternatively, $\mathcal{F}$ can be viewed the subset of Busemann functions $[b]$ so that $b$ restricts to a root on any flat asymptotic to its center. Recall that the star $\operatorname{st}(\tau)$ of a simplex $\tau$ is the union of all Weyl chambers containing it and the Weyl cone $V(o,\operatorname{st}(\tau))$ is the union of points on geodesic rays from $o\in\operatorname{\mathbb{X}}$ to $\operatorname{st}(\tau)$.
Lemma 2.2 (Root pseudometric and Busemann functions).
If $y\in V(x,\operatorname{st}(\tau))$, then
$d_{\alpha}(x,y)=\min\{b(x)-b(y):b\in\mathcal{F}\cap\operatorname{st}(\tau)\}.$
Proof.
Let $e_{1},\dots,e_{d}$ denote the standard basis and $e^{1},\dots,e^{d}$ denote the dual basis. Up to the action of $G$, we may assume that
$x=I_{d},\qquad b(x)-b(y)=-\frac{1}{2}\log\left(\left\lVert e_{1}\right\rVert_{y^{-1}}\left\lVert e^{2}\right\rVert_{y}\right),\quad\text{ and }y=\begin{bmatrix}\lambda^{d-1}&0\\ 0&\lambda^{-1}A\end{bmatrix}$
for a symmetric positive definite matrix $A$ of determinant $1$ with $\sigma_{1}(A)\leq\lambda^{d}$. Then $b(x)-b(y)=\frac{1}{2}\log(\lambda^{d}/A_{11})\geq\frac{1}{2}\log(\lambda^{d}/\sigma_{1}(A))=d_{\alpha}(x,y)$, and equality is achieved when $x,y$ and the center of $b$ lie in a common flat. ∎
2.7. Transversality and parallel sets
A pair $\tau\in\mathbb{KP}^{d-1}$ and $\hat{\tau}\in(\mathbb{KP}^{d-1})^{\ast}$ are called transverse (or antipodal or opposite) if $\tau+\hat{\tau}=\operatorname{\mathbb{K}}^{d}$. When this occurs, there is a parallel set which may be defined by
$P(\hat{\tau},\tau)\coloneqq\{p\in\operatorname{\mathbb{X}}:S_{p}\tau=\hat{\tau}\}.$
Equivalently, if $c$ is a geodesic with $c(+\infty)=\tau$ and $c(-\infty)=\hat{\tau}$ then $P(\hat{\tau},\tau)$ is the union of all points on geodesics parallel to $c$, equivalently, the union of all maximal flats containing the image of $c$. When the points of $\operatorname{\mathbb{X}}$ are interpreted as inner products, the parallel set $P(\hat{\tau},\tau)$ consists of those inner products making $\hat{\tau}$ perpendicular to $\tau$. A parallel set is a totally geodesic submanifold of $\operatorname{\mathbb{X}}$.
3. Angles and distances to parallel sets
In this section we establish certain estimates to be used in the following section. The primary contribution is the angle-to-distance formula in Lemma 3.1.
The following formula generalizes a familiar fact in real hyperbolic geometry. Consider a geodesic with endpoints $x,y$ in the visual boundary and a point $p$. Then the distance from $p$ to the geodesic $xy$ determines the angle at $p$ between $x$ and $y$, and vice versa.
Lemma 3.1 (Angle-to-distance formula).
Let $q\in\operatorname{\mathbb{X}}$, and let $\tau_{+}\in\mathbb{KP}^{d-1}$, $\tau_{-}\in(\mathbb{KP}^{d-1})^{\ast}$ be transverse and let $P(\tau_{-},\tau_{+})$ denote the parallel set. Then
$(d-1)\cos\measuredangle_{q}(\tau_{-},\tau_{+})+d\operatorname{sech}^{2}\left(d(q,P(\tau_{-},\tau_{+})\right)=1.$
Proof.
Let $p\in P=P(\tau_{-},\tau_{+})$ and let $c(t)$ be a geodesic through $p$ normal to $P$. Up to the action of $G=\operatorname{SL}(d,\operatorname{\mathbb{K}})$ and rescaling the speed of $c$, we may assume that $c(t)$ is given by $c(t)=g(t).p$ where
$g(t)^{-1}=\begin{bmatrix}\cosh t&\sinh t&&&\\ \sinh t&\cosh t&&&\\ &&1&&\\ &&&\ddots&\\ &&&&1\\ \end{bmatrix}.$
To compute the angle, we want to use that
$\measuredangle_{g(t)p}(\tau_{-},\tau_{+})=\measuredangle_{p}(g(t)^{-1}\tau_{-},g(t)^{-1}\tau_{+})=\measuredangle_{p}(k_{-}(t)\tau_{-},k_{+}(t)\tau_{+})$
where $g=k_{-}p_{-}$ according to $G=KG_{\tau_{-}}$ and $g=k_{+}p_{+}$ according to $G=KG_{\tau_{+}}$; this is called the Iwasawa or “QR” decomposition. Explicitly, the top $2\times 2$ blocks are given by:
$\frac{1}{\sqrt{\cosh(2t)}}\begin{bmatrix}\cosh t&-\sinh t\\ \sinh t&\cosh t\end{bmatrix},\quad\frac{1}{\sqrt{\cosh(2t)}}\begin{bmatrix}\cosh t&\sinh t\\ -\sinh t&\cosh t\end{bmatrix},$
for $k_{+},k_{-}$ respectively. We have
$k_{+}^{T}k_{-}=\begin{bmatrix}\operatorname{sech}(2t)&\tanh(2t)&&&\\ -\tanh(2t)&\operatorname{sech}(2t)&&&\\ &&1&&\\ &&&\ddots&\\ &&&&1\\ \end{bmatrix},\text{ and }k_{-}^{T}k_{+}=\begin{bmatrix}\operatorname{sech}(2t)&-\tanh(2t)&&&\\ \tanh(2t)&\operatorname{sech}(2t)&&&\\ &&1&&\\ &&&\ddots&\\ &&&&1\\ \end{bmatrix},$
and for $Z=\operatorname{Diag}(d-1,-1,\dots,-1)$ the computation
$\cos\measuredangle_{p}(k_{-}(t)\tau_{-},k_{+}(t)\tau_{+})=\frac{\operatorname{Tr}(k_{-}(-Z)k_{-}^{T}k_{+}Zk_{+}^{T})}{\operatorname{Tr}(Z^{2})}=\frac{1}{d-1}\left(1-d\operatorname{sech}^{2}(2t)\right)$
is straightforward. Our metric yields $d(p,c(t))=d(p,g(t).p)=\sqrt{2\operatorname{Tr}(X^{2})}t=2t$ where $X$ is the unique symmetric matrix such that $g(t)=\exp(tX)$. ∎
Remark 3.2.
Such a formula does not hold for parallel sets in symmetric spaces in general. For example, if $22\leq k\leq d-$ 2, $\tau\in\operatorname{Gr}(k,\operatorname{\mathbb{K}}^{d})$, and $\hat{\tau}\in\operatorname{Gr}(d-k,\operatorname{\mathbb{K}}^{d})$ is transverse to $\tau$, then the distance from $p\in\operatorname{\mathbb{X}}$ to the parallel set $P(\hat{\tau},\tau)$ does not determine the angle $\measuredangle_{p}(\hat{\tau},\tau)$, and vice versa. Similarly, such a formula does not hold for maximal flats in $\operatorname{\mathbb{X}}$ for any $d\geq 3$.
Lemma 3.3 (Detecting transversality).
Let $q\in\operatorname{\mathbb{X}}$, $\tau_{+}\in\mathbb{KP}^{d-1}$ and $\tau_{-}\in(\mathbb{KP}^{d-1})^{\ast}$. Then $\tau_{+}$ is transverse to $\tau_{-}$ if and only if
$\cos\measuredangle_{q}(\tau_{-},\tau_{+})<\frac{1}{d-1}.$
Proof.
If $\tau_{+}$ is transverse to $\tau_{-}$ then this follows from Lemma 3.1.
For the converse, observe that there are only two relative positions for a line and a hyperplane (either the line is in the hyperplane, or the line is transverse to the hyperplane). In fact any Tits angle between ideal points of these fixed types satisfies
$\cos\measuredangle_{\rm Tits}(\tau_{-},\tau_{+})\in\left\{\frac{1}{d-1},-1\right\}.$
Now suppose that $\cos\measuredangle_{q}(\tau_{-},\tau_{+})<\frac{1}{d-1},$ so the Tits angle satisfies between $\tau_{-}$ and $\tau_{+}$ satisfies the same inequality. But then there exists a point $p\in\operatorname{\mathbb{X}}$ whose Riemannian angle between $\tau_{-}$ and $\tau_{+}$ is $\pi$, so they are transverse. ∎
In the next Lemma, we consider a parallel set and an asymptotic geodesic ray. Points far along the ray become exponentially close to the parallel set, with rate depending on the $d_{\alpha}$-pseudometric.
Lemma 3.4 (Distance from ray to parallel set).
Let $p,q\in\operatorname{\mathbb{X}}$ with $d_{\alpha}(x,y)\geq S$ and set $\tau=\zeta(xy)$. Let $\hat{\tau}$ be transverse to $\tau$ and let $P=P(\hat{\tau},\tau)$ be a parallel set with $d(p,P)\leq D$. Then
$d(q,P)\leq\min\{D,(e^{D}-1)e^{-S}\}.$
Proof.
The upper bound of $D$ is immediate from the convexity of the distance function.
Let $r$ be the unique point on $P$ in the horocycle $H(q,\tau)$. Then [Rie21, Lemma 4.10] yields an upper bound for $d(q,P)$ by producing a horocyclic curve $r_{0}$ from $p$ to $r$, pushing it towards $\tau$ to obtain a horocyclic curve $r_{\ell}$ from $q$ to $P$, and estimating its length. The upper bound we use here is a slight modification: the constant $\kappa_{0}$ there becomes $1$ here (due to our normalization of the Riemannian metric), and we integrate the inequality $\left\lvert\dot{r_{\ell}}(t)\right\rvert\leq e^{-S}\left\lvert\dot{r_{0}}(t)\right\rvert\leq e^{t-S}$ over $t$ from 00 to $D$. ∎
Kapovich-Leeb-Porti [KLP14] introduced the $\zeta$-angle which modifies the Riemannian angle by replacing regular directions with a direction in the same Weyl chamber but with type $\zeta$. We need a slight modification of their notion, because we work with a model simplex that is not invariant by the opposition involution $\iota$.
Definition 3.5 ($\zeta$-angle).
Let $x,y,z,w\in\operatorname{\mathbb{X}}$ such that $xy$ is $\iota\zeta$-regular and $xz,xw$ are $\zeta$-regular and let $\hat{\tau}\in(\mathbb{KP}^{d-1})^{\ast}$ and $\tau\in\mathbb{KP}^{d-1}$. Let $\zeta(xz)$ (resp. $\zeta(xw)$) denote the unique ideal point of type $\zeta$ in a common Weyl chamber with $xz(+\infty)$ (resp. $xw(+\infty)$) and let $\iota\zeta(xy)$ denote the unique ideal point of type $\iota\zeta$ in a common Weyl chamber with $xy(+\infty)$. Then set:
$\measuredangle^{\iota\zeta,\zeta}_{x}(y,z)\coloneqq\measuredangle_{x}(\iota\zeta(xy),\zeta(xz)),\quad\measuredangle^{\zeta,\zeta}_{x}(z,w)\coloneqq\measuredangle_{x}(\zeta(xz),\zeta(xw)),$
$\measuredangle^{\iota\zeta,\zeta}_{x}(\hat{\tau},z)\coloneqq\measuredangle_{x}(\hat{\tau},\zeta(xz)),\quad\measuredangle^{\iota\zeta,\zeta}_{x}(y,\tau)\coloneqq\measuredangle_{x}(\iota\zeta(xy),\tau).$
We conclude this section with an estimate controlling $\zeta$-angles between $x,y\in\operatorname{\mathbb{X}}$ as seen from $o\in\operatorname{\mathbb{X}}$ which is useful when the pseudodistances $d_{\alpha}(o,x),d_{\alpha}(o,y)$ are larger than $d(x,y)$.
Lemma 3.6 ($\zeta$-angle estimate).
Let $o\in\operatorname{\mathbb{X}}$ and $S>D\geq 0$. The map
$f=\zeta(o\cdot)\colon\{x\in\operatorname{\mathbb{X}}:d_{\alpha}(o,x)>0\}\to(K\cdot Z,\measuredangle)$
is smooth, and for $x,y\in\operatorname{\mathbb{X}}$ such that $d_{\alpha}(o,x)\geq S$ and $d(x,y)\leq D$ we have
$\measuredangle_{o}(\zeta(ox),\zeta(oy))\leq\frac{\alpha(Z)}{\sinh(S-D)}d(x,y).$
Proof.
Observe that the open Weyl cone
$V=\left\{q=\begin{bmatrix}\lambda&0\\ 0&A\end{bmatrix}:\lambda>\sigma_{1}(A)\right\}$
is the preimage of $Z$ under $f$. Since $V$ is an open subset of a parallel set, it is a smooth submanifold of $\operatorname{\mathbb{X}}$. Let
$\mathfrak{k}^{Z}=\left\{U=\begin{bmatrix}0&-u^{\dagger}\\ u&0\end{bmatrix}:u\in\mathbb{K}^{d-1}\right\}.$
It is easy to check that for all $q$ in $V$ and $U\in\mathfrak{k}^{Z}$, the fundamental vector field $U^{\ast}_{q}$ is orthogonal to $T_{q}V$. Mapping $K\times V$ by $(k,v)\mapsto kv$ descends to a $K$-equivariant diffeomorphism between $K\times_{K_{Z}}V$ and $X=\{q\in\operatorname{\mathbb{X}}:d_{\alpha}(o,q)>0\}$. It follows that $f$ is smooth.
We use that for a $C^{1}$ function between Riemannian manifolds, $f\colon X\to K\cdot Z$, the optimal Lipschitz constant is given by $\sup\left\{\left\lvert df_{q}\right\rvert:q\in X\right\}$. Since the action of $K$ on $K\cdot Z$ is transitive and $f$ is equivariant, for any $v\in T_{q}X$ there exists $U\in\mathfrak{k}^{f(q)}$ such that $df_{q}(v)=U^{\ast}_{f(q)}=df_{q}(U^{\ast}_{q})$. It is convenient to introduce the alternative notation $\mathrm{ev}_{q}(U)=U^{\ast}_{q}$ for the fundamental vector field (read “evaluated at $q$”). Since the decomposition $T_{q}X=\ker(df_{q})\oplus\mathrm{ev}_{q}(\mathfrak{k}^{f(q)})$ is orthogonal, we have $\left\lvert v\right\rvert\geq\left\lvert U^{\ast}_{q}\right\rvert$, so
$\frac{\left\lvert df_{q}(v)\right\rvert}{\left\lvert v\right\rvert}\leq\frac{\left\lvert U^{\ast}_{f(q)}\right\rvert}{\left\lvert U^{\ast}_{q}\right\rvert}.$
For a matrix $B$ let $\left\lVert B\right\rVert^{2}=2\operatorname{Tr}(B^{\dagger}B)$. For $U\in\mathfrak{k}^{Z}$ write $U=\sum_{\beta}U_{\beta}$ for its root space decomposition; this agrees with its decomposition into matrix entries. We compute the norm squared of the vector $U^{\ast}_{Z}$ in the Euclidean space of Hermitian matrices:
$\left\lVert U^{\ast}_{Z}\right\rVert^{2}=\left\lVert[Z,U]\right\rVert^{2}=\left\lVert\sum_{\beta(Z)\neq 0}\beta(Z)U_{\beta}\right\rVert^{2}=\alpha(Z)^{2}\left\lVert U\right\rVert^{2}.$
For $q\in V$ satisfying $d_{\alpha}(o,q)\geq S^{\prime}$ we compute the norm squared of the vector $U^{\ast}_{Z}$ in $T_{q}\operatorname{\mathbb{X}}$ with respect to the Riemannian metric:
$\left\lvert U^{\ast}_{q}\right\rvert^{2}_{q}=\left\lvert\mathrm{ev}_{o}\circ\operatorname{Ad}(q)^{-1/2}(U)\right\rvert^{2}_{o}=\left\lvert\sum_{\beta(Z)>0}\left(e^{\beta(\vec{d}(o,q))}-e^{-\beta(\vec{d}(o,q))}\right)(U_{\beta})^{\ast}_{o}\right\rvert^{2}_{o}\geq\sinh^{2}(S^{\prime})\left\lVert U\right\rVert^{2}.$
To conclude we observe that the geodesic segment from $x$ to $y$ in $\operatorname{\mathbb{X}}$ lies in $\{q\in\operatorname{\mathbb{X}}:d_{\alpha}(o,q)\geq S-D\}$. ∎
4. Straight and spaced sequences
Kapovich-Leeb-Porti [KLP17, KLP18a] and Bochi-Potrie-Sambarino [BPS19] have proven that a finitely generated subgroup $\Gamma$ of $\operatorname{SL}(d,\operatorname{\mathbb{K}})$ is projective Anosov if and only if every geodesic in $\Gamma$ maps to a uniformly $d_{\alpha}$-undistorted sequence in $\operatorname{\mathbb{X}}$, see Definition 4.1. In this section we state and prove a local criterion for a sequence in $\operatorname{\mathbb{X}}$ to be globally and uniformly $d_{\alpha}$-undistorted.
Theorem 4.2 is similar to [KLP23, Theorem 3.18] and [Rie21, Theorem 5.1] but subtly different. The statement here assumes that the sequence is $S$-spaced with respect to the $d_{\alpha}$-pseudometric, which is a bit weaker than assuming that consecutive pairs are simultaneously uniformly regular and $S$-spaced with respect to the Riemannian metric. The conclusion is also weaker, since we only obtain that the sequence is $d_{\alpha}$-undistorted, and it may fail to fellow travel truncated Weyl cones. However, once this statement, which is purely about sequences in the symmetric space, is applied to actions of finitely generated subgroups where the sequnces come from geodesics in the Cayley graph, the Lipschitz property of the orbit map implies the sequences are Morse. It is straightforward to modify the proof of Theorem 4.2 to obtain a statement which assumes this stronger condition and implies that the sequence is uniformly Morse.
We need the following definitions in order to state the main theorem.
Definition 4.1.
Let $(x_{n})$ be a sequence of points in $\operatorname{\mathbb{X}}$.
  1. (1)
    The sequence is $d_{\alpha}$-undistorted with constants $c_{1}>0$ and $c_{2}\geq 0$ if for all $m,n$:
    $c_{1}\left\lvert m-n\right\rvert-c_{2}\leq d_{\alpha}(x_{n},x_{m}).$
  2. (2)
    The sequence is $S$-spaced for $s\geq 0$ if for all $n$:
    $d_{\alpha}(x_{n},x_{n+1})\geq S.$
  3. (3)
    The sequence is $\epsilon$-straight if each segment $x_{n}x_{n+1}$ is $\zeta$-regular and for all $n$:
    $\measuredangle^{\iota\zeta,\zeta}_{x_{n}}(x_{n-1},x_{n+1})\geq\pi-\epsilon.$
We may now state and prove the main theorem: sufficiently straight and spaced sequences are $d_{\alpha}$-undistorted.
Theorem 4.2 (Sufficiently straight and spaced sequences are $d_{\alpha}$-undistorted).
Let $\alpha_{new}<\alpha_{0}$, $\epsilon_{aux}>\epsilon$, $S$ and $\delta_{1},\delta_{2},\delta_{3}$ satisfy:
  1. (1)
    $\min\left\{\cos\left(\epsilon_{aux}+\frac{\delta_{4}\zeta_{0}}{\sinh(S-\delta_{4})}\right),\cos\left(2\epsilon_{aux}-\epsilon\right)\right\}>-\frac{1}{d-1}$
    so that certain simplices are antipodal by Lemma 3.3,
  2. (2)
    $\max\left\{(1-d)\cos(\epsilon_{aux})+d\operatorname{sech}^{2}(\delta_{1}),(1-d)\cos(2\epsilon_{aux}-\epsilon)+d\operatorname{sech}^{2}(\delta_{3})\right\}\leq 1$
    so that certain points are close to parallel sets by Lemma 3.1,
  3. (3)
    $(1-d)\cos(\epsilon_{aux}-\epsilon)+d\operatorname{sech}^{2}(\delta_{2})\leq 1$
    so that certain $\zeta$-angles are small enough by Lemma 3.1,
  4. (4)
    $(e^{\delta_{1}}-1)e^{-S}\leq\delta_{2}$
    so that certain points are close to parallel sets by Lemma 3.4, and
  5. (5)
    $\delta_{4}\geq\min\{2\delta_{3},\delta_{3}+(e^{\delta_{3}}-1)e^{-S}\}\text{ and }S>2\delta_{4}$
    so that certain projections to Weyl cones are uniformly monotonic.
Then an $S$-spaced and $\epsilon$-straight sequence is $d_{\alpha}$-undistorted with constants $(S-2\delta_{4},2\delta_{4})$.
Remark 4.3.
The reader may want to convince themself that for $\epsilon$ sufficiently small and $S$ sufficiently large, there exist auxiliary parameters satisfying the hypotheses of Theorem 4.2. To see this, choose any $\epsilon$ smaller than $\epsilon_{max}=\cos^{-1}\left(\frac{-1}{d-1}\right)$ and any $\epsilon_{aux}$ satisfying $\epsilon<\epsilon_{aux}<\frac{\epsilon+\epsilon_{max}}{2}$. Then for any $\delta_{1},\delta_{2},\delta_{3}$ satisfying Assumptions 2 and 3, one observes that for $S$ sufficiently large, Assumptions 1, 4 and 5 are satisfied.
Proof.
Step 1: Propagation. We first need to show that under these assumptions, the property of “moving $\epsilon_{aux}$-away from/towards a hyperplane/line” propagates along the sequence. Assume that $\hat{\tau}\in(\mathbb{KP}^{d-1})^{\ast}$ satisfies
$\measuredangle^{\iota\zeta,\zeta}_{x_{0}}(\hat{\tau},x_{1})\geq\pi-\epsilon_{aux}.$
Write $\tau_{01}\in\mathbb{KP}^{d-1}$ for $\zeta(x_{0}x_{1}(+\infty))$. By Assumption 1 and Lemma 3.3, we have that $\hat{\tau}$ is transverse to $\tau_{01}$. By Assumption 2 and Lemma 3.1, we have
$d(x_{0},P(\hat{\tau},\tau_{01}))\leq\operatorname{sech}^{-1}\left(\sqrt{\frac{1}{d}\left(1-(1-d)\cos(\epsilon_{aux})\right)}\right)\leq\delta_{1}.$
By Assumption 4 and Lemma 3.4, we have
$d(x_{1},P(\hat{\tau},\tau_{01}))\leq(e^{\delta_{1}}-1)e^{-\alpha_{0}s}\leq\delta_{2}$
small enough that Assumption 3 and Lemma 3.1 imply
$\measuredangle^{\iota\zeta,\iota\zeta}_{x_{1}}(\hat{\tau},x_{0})\leq\pi-\cos^{-1}\left(\frac{1}{d-1}\left(1-d\operatorname{sech}^{2}\left(\delta_{2}\right)\right)\right)\leq\epsilon_{aux}-\epsilon.$
By straightness and the triangle inequality, we have
$\measuredangle^{\iota\zeta,\zeta}_{x_{1}}(\hat{\tau},x_{2})\geq\pi-\epsilon_{aux}.$
By induction, we have that $\measuredangle^{\iota\zeta,\zeta}_{x_{n}}(\hat{\tau},x_{n+1})\geq\pi-\epsilon_{aux}$ for all $n\geq 0$. A similar proof shows that the property of moving $\epsilon_{aux}$-towards a line propagates along the sequence.
Step 2: Extraction. The previous step allows us to find simplices $\tau_{\pm}$ that the sequence moves $\epsilon_{aux}$-towards/away from. Indeed, let
$C^{+}_{n}=\{\tau_{+}:\measuredangle^{\iota\zeta,\zeta}_{x_{n}}(x_{n-1},\tau_{+})\geq\pi-\epsilon_{aux}\}\text{ and }C^{-}_{n}=\{\tau_{-}:\measuredangle^{\iota\zeta,\zeta}_{x_{n}}(\tau_{-},x_{n+1})\geq\pi-\epsilon_{aux}\}.$
The previous step implies that $\bigcap_{n}C^{\pm}_{n}$ is nonempty, so we may extract $\tau_{\pm}\in\bigcap_{n}C^{\pm}_{n}$. The proof of the previous step shows that since $\tau_{-}\in C^{-}_{n-1}$ we have
$\measuredangle^{\iota\zeta,\zeta}_{x_{n-1}}(\tau_{-},x_{n})\geq\pi-\epsilon_{aux}\implies\measuredangle^{\iota\zeta,\iota\zeta}_{x_{n}}(\tau_{-},x_{n-1})\leq\epsilon_{aux}-\epsilon$
and a similar statement holds for $\tau_{+}$, so the triangle inequality gives
$\left\lvert\measuredangle^{\iota\zeta,\zeta}_{x_{n}}(\tau_{-},\tau_{+})-\measuredangle^{\iota\zeta,\zeta}_{x_{n}}(x_{n-1},x_{n+1})\right\rvert\leq\measuredangle^{\iota\zeta,\iota\zeta}_{x_{n}}(\tau_{-},x_{n-1})+\measuredangle^{\zeta,\zeta}_{x_{n}}(\tau_{+},x_{n+1})\leq 2(\epsilon_{aux}-\epsilon).$
By straightness, the previous inequality implies that $\measuredangle^{\iota\zeta,\zeta}_{x_{n}}(\tau_{-},\tau_{+})\geq\pi-2\epsilon_{aux}+\epsilon$, so $\tau_{-}$ is transverse to $\tau_{+}$ by Assumption 1 and Lemma 3.3. Moreover,
$d(x_{n},P(\tau_{-},\tau_{+}))\leq\operatorname{sech}^{-1}\left(\sqrt{\frac{1}{d}\left(1-(1-d)\cos(2\epsilon_{aux}-\epsilon)\right)}\right)\leq\delta_{3}$
by Assumption 2 and Lemma 3.1.
Step 3: Undistortion. We now verify that the sequence $(x_{n})$ is $d_{\alpha}$-undistorted. For all $n$, let $\overline{x}_{n}$ denote the nearest point to $x_{n}$ in the parallel set $P(\tau_{-},\tau_{+})$. We will show that the sequence $(\overline{x}_{n})$ is $d_{\alpha}$-undistorted.
Let $\xi$ be the ideal point corresponding to the ray $\overline{x}_{n}\overline{x}_{n+1}$. Since the rays $x_{n}\xi$ and $\overline{x}_{n}\xi$ are asymptotic, their Hausdorff distance is at most $d(x_{n},\overline{x}_{n})\leq\delta_{3}$, so, by the proof of Lemma 3.4, $x_{n+1}$ is at most $\delta_{4}$ from $x_{n}\xi$. We then have
$\measuredangle_{\rm Tits}^{\iota\zeta,\zeta}(\tau_{-},\xi)\geq\measuredangle_{x_{n}}^{\iota\zeta,\zeta}(\tau_{-},\xi)\geq\measuredangle_{x_{n}}^{\iota\zeta,\zeta}(\tau_{-},x_{n+1})-\measuredangle_{x_{n}}^{\zeta,\zeta}(x_{n+1},\xi)\geq\pi-\epsilon_{aux}-\measuredangle_{x_{n}}^{\zeta,\zeta}(x_{n+1},\xi)$
and we can bound
$\measuredangle^{\zeta,\zeta}_{x_{n}}(x_{n+1},\xi)\leq\frac{\delta_{4}\zeta_{0}}{\sinh(S-\delta_{4})}$
by Lemma 3.6.
By Assumption 1, it follows that $\tau_{-}$ is antipodal to $\zeta(\xi)$; since $\tau_{+}$ is the only simplex in $\partial_{\rm vis}P(\tau_{-},\tau_{+})$ antipodal to $\tau_{-}$, this implies $\tau(\xi)=\tau_{+}$. By the convexity of Weyl cones the sequence of projections $\overline{x}_{n}$ land in nested Weyl cones in $P(\tau_{-},\tau_{+})$ [KLP17, Corollary 2.11].
Finally we show that $(\overline{x}_{n})$ is $d_{\alpha}$-undistorted. Note that the vector-valued triangle inequality (see [KLP17, Par] or [Rie21, Corollary 3.8]) implies that $d_{\alpha}(\overline{x}_{n},\overline{x}_{n+1})\geq d_{\alpha}(x_{n},x_{n+1})-d(x_{n},\overline{x}_{n})-d(x_{n+1},\overline{x}_{n+1})\geq S-2\delta_{4}$, which is positive by Assumption 5. Fix $m>n$. By Lemma 2.2 and the nestedness of Weyl cones, there exists a Busemann function $b$ so that
$d_{\alpha}(\overline{x}_{n},\overline{x}_{m})$ $=b(\overline{x}_{n})-b(\overline{x}_{m})=b(\overline{x}_{n})-b(\overline{x}_{n+1})+b(\overline{x}_{n+1})-b(\overline{x}_{n+2})+\cdots+b(\overline{x}_{m-1})-b(\overline{x}_{m})$
$\geq d_{\alpha}(\overline{x}_{n},\overline{x}_{n+1})+d_{\alpha}(\overline{x}_{n+1},\overline{x}_{n+2})+\cdots+d_{\alpha}(\overline{x}_{m-1},\overline{x}_{m})\geq(m-n)(S-2\delta_{4}).$
This implies that $(x_{n})$ is $d_{\alpha}$-undistorted with constants $(S-2\delta_{4},2\delta_{4})$. ∎
5. Example
We illustrate the practicality of the algorithm with an example. We consider a fixed surface subgroup of $\operatorname{SL}(3,\operatorname{\mathbb{R}})$ and verify the Anosov property via a computation involving only the words of length $8$. Using KBMAG [HGT23], the words of length $8$ can be enumerated via a finite state automoton. Then via [Rie24] or [Wei24], the straightness and spacing parameters of the associated midpoint sequence can be computed. It is then easy to obtain auxiliary parameters satisfying Theorem 4.2, which certifies the Anosov condition.
Let $\Gamma$ be the subgroup of $\operatorname{SL}(3,\operatorname{\mathbb{R}})$ generated by
$S=\left\{\begin{bmatrix}\cos(\theta)&\sin(\theta)&0\\ -\sin(\theta)&\cos(\theta)&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}\cosh(T)&0&\sinh(T)\\ 0&1&0\\ \sinh(T)&0&\cosh(T)\end{bmatrix}\begin{bmatrix}\cos(\theta)&-\sin(\theta)&0\\ \sin(\theta)&\cos(\theta)&0\\ 0&0&1\end{bmatrix}:\theta\in\left\{0,\frac{\pi}{8}.\frac{\pi}{4},\frac{3\pi}{8}\right\}\right\}$
where $T=2\cosh^{-1}(\cot(\pi/8))$.
Labelling the generators by $a,b,c,d$ and their inverses by $A,B,C,D$, a presentation of $\Gamma$ is given by the single relation $adCbADcB$. The words of length at most $8$ in $\Gamma$ can be enumerated using the KBMAG package in GAP [HGT23]. For each such word $w$, we decompose it as $w=w_{1}w_{2}$ into words of length $4$.
We set
$m_{1}=\operatorname{mid}(o,w_{1}^{-1}o),\text{ and }m_{2}=\operatorname{mid}(o,w_{2}o)$
and compute
$s_{w}=d_{\alpha}(m_{1},m_{2}),\quad\epsilon_{w}^{+}=\measuredangle_{m_{1}}^{\zeta,\zeta}(o,m_{2}),\text{ and }\epsilon_{w}^{-}=\measuredangle_{m_{2}}^{\iota\zeta,\iota\zeta}(o,m_{1}).$
We now consider a geodesic $(\gamma_{n})$ in $\Gamma$ with $\gamma_{0}=id$. We will see that $\gamma_{n}o$ is $d_{\alpha}$-undistorted sequence in $\operatorname{\mathbb{X}}$, equivalently, that $\gamma_{4n}o$ is $d_{\alpha}$-undistorted, equivalently, that the sequence of midpoints $m_{n}=\operatorname{mid}(\gamma_{4n}o,\gamma_{4n+4}o)$ is $d_{\alpha}$-undistorted. In general, if a finitely generated subgroup $\Gamma$ is Anosov then for some $k>0$ the sequence of midpoints $m_{n}=\operatorname{mid}(\gamma_{kn}o,\gamma_{kn+k}o)$ is straight-and-spaced, which can be verified by examining words of length $3k$, see [KLP14, Proposition 3.32] or [Rie21, Theorem 5.4]. We use the same trick as [Rie23] to instead consider words of length $2k$, where $k=4$ in this example. The idea is that one can estimate the straightness parameter for words of length $3k$ with words of length $2k$.
Since $\Gamma$ acts by isometries, the sequence of midpoints is $S$-spaced for
$S=\min\{s_{w}:\left\lvert w\right\rvert=8\},$
and $\epsilon$-straight for
$\epsilon=\max\{\epsilon_{w}^{+}:\left\lvert w\right\rvert=8\}+\max\{\epsilon_{w}^{-}:\left\lvert w\right\rvert=8\}.$
We compute these in the example in the code available at [Rie24], obtaining
$\min\{\cos(\epsilon_{w}^{+}):\left\lvert w\right\rvert=8\},\min\{\cos(\epsilon_{w}^{-}):\left\lvert w\right\rvert=8\}\approx 0.87\implies\epsilon\approx 2\cos^{-1}(0.87)\approx 1.03,$
$S\approx 3.08.$
We choose auxiliary constants $\epsilon_{aux}=0.7\epsilon+0.3\epsilon_{max}$, where $\epsilon_{max}=\cos^{-1}\left(\frac{-1}{3-1}\right)$, and
$\delta_{1}=\operatorname{sech}^{-1}\left(\sqrt{\frac{1}{3}(1-(1-3)\cos(\epsilon_{aux}))}\right)\approx 0.92$
$\delta_{2}=\operatorname{sech}^{-1}\left(\sqrt{\frac{1}{3}(1-(1-3)\cos(\epsilon_{aux}-\epsilon))}\right)\approx 0.18$
$\delta_{3}=\operatorname{sech}^{-1}\left(\sqrt{\frac{1}{3}(1-(1-3)\cos(2\epsilon_{aux}-\epsilon))}\right)\approx 1.29$
$\delta_{4}=\delta_{3}+(e^{\delta_{3}}-1)e^{-S}\approx 1.41.$
These constants satisfy Theorem 4.2. So each sequence of midpoints $(m_{n})$ constructed above is $d_{\alpha}$-undistorted. This implies that every geodesic in $\Gamma$ maps to a $d_{\alpha}$-undistorted sequence in $\operatorname{\mathbb{X}}$. By [KLP18b] or [BPS19], it follows that $\Gamma$ is Anosov.